The members of Peace Club decided to raise Php4000 to help in shouldering the cost of peace advocacy campaigns. The amount to be raised was divided equally among the members. Due to increase of 20 members , the amount to be raised by each member went down by Php10. How many members were originally in the club?
Answer by jorel1380(3719)

Let n be the amount of members in the Peace club. Then, n+20 equals the amount of members plus new members. So:
4000/n=4000/(n+20) +10
4000(n+20)=4000n+10(nē+20n)
4000n+80000=10nē+4200n
10nē+200n-80000=0
nē+20n-8000=0
solve for n
☺☺☺☺